OpenStudy (anonymous):

Which expression describes the rule of this sequence? Three, six, nine, twelve, fifteen, eighteen, and so on.

OpenStudy (anonymous):

hold on a sec

OpenStudy (anonymous):

k :)

OpenStudy (anonymous):

wait wrong question

OpenStudy (anonymous):

Which expression describes the rule of this sequence? Three, six, nine, twelve, fifteen, eighteen, and so on. three plus the quantity n minus one three times n three plus the quantity three times n three plus n

OpenStudy (anonymous):

Aris Brain (5,781) It depends of what n is defined to be. If n is the set of integers 1, 2, 3, 4, and so on to infinity then the expression that describes the sequence is 3 times n. If n is defined as the set 0, 3, 6, 9, etc. , then the expression that describes the sequence is n plus 3 from webanswers.com

OpenStudy (anonymous):

ignore the first part

OpenStudy (anonymous):

lol chu got that from online

OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence? three, seven, eleven four, eight, twelve three, six, nine four, seven, eleven

OpenStudy (anonymous):

@Purplerainbowcherry

OpenStudy (anonymous):

4n-1 , so if n=1, you would get 4-1 = 3 if n=2 , you would get 4*2 -1 = 8-1 = 7 so i would choose the first choice

OpenStudy (anonymous):

Your teacher has a jar of chocolate candies. For every completed assessment, she gives herself four pieces of chocolate. Part a. On your own paper, complete the table below to show the relationship between the number of assessments and pieces of chocolate eaten. Table with 16 columns, first column has Number of Assessments, second column has Pieces of Chocolate the other 14 columns are empty Part b. Using the values in your table, look for the relationship between the number of completed assessments and number of chocolate candies eaten. Use this relationship to write an expression that represents the number of assessments completed and the number of candies eaten. (3 points) Part c. Use your expression to show the number of candies eaten after 67 completed assessments. Show all of your work. (3 points)

OpenStudy (anonymous):

I need this in essay form

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Purplerainbowcherry